A person has purchased 10 of 1000 tickets sold in raffle. To determine five prize winners, 5 tickets were drawn at random without replacement. Compute the probabilty that this person wins at least one prize.
Answer by richwmiller(17219)

10/1000=1/100
5 *1/100=5/100=1/20